The normalized wave function for a hydrogen atom in the 1s state is given by
ψ(r) =  ας παοwhere α0 is the Bohr radius, which is equal to 5.29 × 10-11 m. What is the probability of finding the electron at a distance greater than 7.8 α0 from the proton?

Anwer is 2.3 × 10-5, but how can I get it?
Answer #1

probability distribution punction P = 4pi r^2* |ψ(r)|^2 = 4r^2/[a0^3] e^(-2r/a0)

probability = integral Pdr

=  integral 4r^2/[a0^3] e^(-2r/a0) dr    limit from 7.8a0 to infinity

= [a0^2+2a0x+2x^2]*-a0/[a0^3]*e^(-2r/a0) from r=7.8a0 to infinity

= [1+2*7.8 +2*7.8^2]* e^(-15.6)

= 2.3*10^-5 Answer